# Large Language Models in Qualitative Research: Can We Do the Data Justice?

### v1

@@ 4.3.5. Who is in charge or creating meaning from human experience?
Many researchers in our participant group discussed how qualitative methods are well-suited to understanding the why behind human experience, (P1, P2, P6, P10, P13, P16, P20), particularly in contrast to what quantitative approaches can tell us. One theme is resounding: qualitative researchers want to keep humans in charge and center their perspectives. Summarizing this point, P8 shared, “The whole point of doing qualitative research is to engage with people, and if we just stop doing that, then I feel like it endangers qualitative research in some way.” P15 took the view using LLMs in qualitative research may still be compatible with this goal as long as humans are the ones ultimately determining meaning from an analysis: ”I think LLMs would be incredible for topical emergence. Only humans can really [tell] you what the themes are ultimately.” Can LLMs surface useful complementary insight while we keep humans in charge of what those features mean in the context of human lives? LLMs do not have lived experience, and do not “understand” the lived experiences of humans. In a set of methodologies concerned with making sense of the “texture of human lives” (P19), keeping the project of creating meaning as a human-centered project is critical.
@@ 5. Discussion
Key themes emerged from interviews that allow us to contextualize the current moment, and imagine how qualitative researchers may adopt LLMs moving forward while staying faithful to the responsibilities they have towards participants. The qualitative researchers we interviewed were largely open to or interested in the conscientious use of LLMs across the research pipeline. Previous work has evaluated specific systems for qualitative researchers that leverage LLMs or other AI (Gao et al., 2023a; Gebreegziabher et al., 2023; Overney et al., 2024), finding promising use cases for qualitative researchers. In contrast, we studied participants’ engagement with LLMs as they themselves had chosen, independent of any one tool or prescribed use case. In participant answers, we found optimism about many potential uses but also valid concerns about how to align these uses to their personal and disciplinary research values. We found this across participants at many stages of familiarization (Long et al., 2024), between arization (not familiar with LLM tools), familiarization (actively playing with LLM tools), and post-familiarization (deploying LLMs for specific chosen use cases).
A second major finding is that although researchers are increasingly surrounded by task-focused AI tools, their imagination of “AI” or “LLMs” is dominated by the chatbot interaction pattern exemplified by ChatGPT. Participants are comfortable with task-specific tools that use LLMs such as grammar check, speech-to-text transcription, and translation, and use them extensively with generally (but not universally) high levels of trust. Their perception of the suitability of LLMs for research, however, is largely focused on the single ChatGPT system and the applications easiest for it, such as ideation and text editing; even among this more computational group, only the most technologically savvy participants were currently leveraging LLMs to do more complex tasks such as rigorous qualitative coding.
ChatGPT and its competitors are specifically designed to emulate a helpful assistant. For participants, this design decision increasingly blurs the boundary between seeing LLMs as a “tool” or as a “collaborator.” In contrast to previous contested technological advances in QDA software tools, ChatGPT is uniquely compelling because it performs human-like interaction patterns such as following natural language instructions, and gives the appearance of recognizing nuance in text described in abstract terms, not just counting words or providing statistics. Prior research supports this finding: the chatbot is an example of a user interface metaphor (Neale and Carroll, 1997); specific systems drive perceptions of broader technologies with Galloway’s interface effect (Galloway, 2012); and brief interactions with systems produce misleading impressions of overall performance (Bowman, [n. d.]). As these effects are unlikely to be limited to qualitative researchers, this work has potential impact in many domains beyond HCI.
@@
On top of surfacing researchers’ points of excitement for future uses of LLMs to assist their research, we outline the many concerns they have, including uncertainties about norms, privacy, and output quality that could diminish the research process or misrepresent participants. Worries about norms and guidelines are ubiquitous, despite the active development and publication of norms by institutions and publication venues (e.g. ACM policies), ethical and supervisory boards, and academic research papers (Watkins, 2023; Hosseini and Horbach, 2023). In the social norm change literature, scholars posit that there is often a tipping point where social norms change rapidly (Andrighetto and Vriens, 2022). Our research suggests that some norms regarding LLMs may be close to a clear tipping point, such as the understanding that identifiable participant data ought not be sent to a third-party tool that may train AI models on the data, while norms regarding some other concerns have so far eluded convergence. Our work confirms prior perspectives that AI evolution is outpacing our norm formation processes (Baronchelli, 2024; Feldstein, 2024), and we show how scholars feel pressure and uncertainty when justifying LLM-aided techniques for academic publications. We show these systems being adopted despite wide-ranging concerns about privacy, suggesting a structural failure to educate researchers on best practices—a concern likely implicating research across fields, agnostic to methodology.
@@
Beyond these practical concerns, a set of fundamental tensions arose in considering the use of LLMs to conduct qualitative research. There has long been a tension between work seeking to combine computational power with interpretive approaches (Baumer et al., 2017; Nelson, 2020) and work arguing that qualitative values must be protected from quantitative evaluation standards (Crabtree, 2024; Soden et al., 2024). Our interviews suggest that LLMs may continue blurring the line between positivist and interpretive approaches, as exemplified in emergent work (Lam et al., 2024; Choksi et al., 2024). As Messeri and Crockett (2024) discuss, the way some approaches are more readily aided by LLMs may change the research landscape by making some questions more tractable than others. There is a danger that the ability to analyze data at scale with less human involvement may make situated, slow, contextual and interpretivist approaches to qualitative data less attractive because of quick publication cycles and the perceived objectivity of LLMs. Mixed method researchers, who make up roughly half of our participants, may be positioned to accelerate this change. Published work uses LLMs for many types of text analysis, including in powerful new tools that are increasingly positioned to facilitate “theory-driven analysis” of textual data (Lam et al., 2024) and they may be adopted by users regardless of methodology. As LLMs produce summaries, topics, and labels of textual data that look increasingly fluent, nuanced, and contextual, researchers may increasingly be tempted to use these outputs in analysis while decreasing the type of engagement with participants and their data that researchers say is central to qualitative research values.

### v2

@@ 4.3.5. Who is in charge or creating meaning from human experience?
Many researchers in our participant group discussed how qualitative methods are well-suited to understanding the why behind human experience, (P1, P2, P6, P10, P13, P16, P20), particularly in contrast to what quantitative approaches can tell us. One theme is resounding: qualitative researchers want to keep humans in charge and center their perspectives. Summarizing this point, P8 shared, “The whole point of doing qualitative research is to engage with people, and if we just stop doing that, then I feel like it endangers qualitative research in some way.” P15 took the view using LLMs in qualitative research may still be compatible with this goal as long as humans are the ones ultimately determining meaning from an analysis: ”I think LLMs would be incredible for topical emergence. Only humans can really [tell] you what the themes are, ultimately.”
@@ 5. Discussion
Our interviews situate at least a partial answer on how qualitative researchers can leverage LLMs while staying faithful to responsible, participant-centered research. In response to RQ1, we find that many qualitative researchers are open to the conscientious use of LLMs across the research pipeline. RQ2 poses the question of which tensions might arise if they are used, and we find that qualitative researchers harbor many valid questions and concerns about how to align LLM uses to their personal and disciplinary research values. RQ3 raises the idea of potential models for ethical use of LLMs in qualitative research, which we discuss in Section 6.
5.1. From Task-Specific to Ubiquitous AI
We build on prior work on human-AI collaboration in qualitative contexts (Jiang et al., 2021; Feuston and Brubaker, 2021) by situating AI use in the current moment. Feuston and Brubaker (2021) argued that successful human-AI collaboration for qualitative research requires the intentional use of AI for specific tasks, such that researchers are not “displaced as the primary analysts.”  Jiang et al. (2021) similarly suggest future research explore which precise sub-tasks of qualitative research are most appropriate for AI involvement. Additionally, many papers on human-AI collaborations reported that the order of AI involvement matters in qualitative research, since researchers do not want to be influenced by suggestions too early in their bottom-up sensemaking process  (Lam et al., 2024; Overney et al., 2024; Jiang et al., 2021; Feuston and Brubaker, 2021). Aligning with this idea, our participants were comfortable with task-specific tools that use LLMs such as grammar check, speech-to-text transcription, and translation, and used them extensively.
However, our findings highlight that the ideal of intentionally using specific AI tools for clearly delineated research tasks is increasingly challenging to meet, due to the prevalence of AI through flexible, seemingly all-purpose chatbot interaction patterns popularized by ChatGPT. Participants used AI across their research pipeline, a finding also supported by  Liao et al. (2024). A natural consequence of this shift is that the lines between AI and non-AI tools blur, making intentional use more challenging.  Narayanan and Kapoor (2024) speculate that generative AI will become “part of our digital infrastructure, instead of being a tool people use for specific purposes.” Our findings suggest that this shift is underway.
Nonetheless, this shift need not be solely negative. Our participants indicated that the “helpful assistant” nature of the most popular AI tools today collapse the boundary between seeing LLMs as a “tool” and as a “collaborator.” While there are significant risks of overly anthropomorphizing AI (Messeri and Crockett, 2024), these natural language interactions may also be the kind of human-AI collaboration that “support serendipity,” which  Jiang et al. (2021) champion.
5.2. The Sociotechnical Breakdown of Guidelines, Norms, and Tooling
@@
Worries about adhering to norms and guidelines are ubiquitous, despite the active development and publication of norms by institutions and publication venues (e.g. ACM policies), ethical and supervisory boards, and academic research papers (Watkins, 2023; Hosseini and Horbach, 2023). In the social norm change literature, scholars posit that there is often a tipping point where social norms change rapidly (Andrighetto and Vriens, 2022). Our research suggests that some norms regarding LLMs may be close to a clear tipping point, such as the understanding that identifiable participant data ought not be sent to a third-party tool that may train AI models on the data, while norms regarding some other concerns have so far eluded convergence. Our work confirms prior perspectives that AI evolution is outpacing our norm formation processes (Baronchelli, 2024; Feldstein, 2024), and we show that scholars feel pressure and uncertainty when justifying LLM-aided techniques for academic publications.
We show that LLMs are being broadly adopted in qualitative research practices despite wide-ranging concerns about privacy. Ongoing work suggests that non-computing researchers have more pronounced ethical concerns in using LLMs (Liao et al., 2024), and our findings suggest that ethics and privacy were top of mind for qualitative researchers more broadly. Nonetheless, the present research suggests that participants use or consider using LLMs despite these concerns, and sometimes use these tools while still being unsure about appropriate use.  Kapania et al. (2024) similarly find that researchers are often aware of potential ethical issues, but uncertain about how to address them. We position the current research environment as a fundamental sociotechnical breakdown: the available tools do not present researchers with the right options to confidently preserve privacy, the necessary guidelines and norms are lagging behind, and the social context is increasingly normalizing and even expecting the use of AI tools. This misalignment of incentives resembles Ackerman’s notion of the socio-technical gap (Ackerman, 2000), which refers to the gap between technical capabilities and social expectations.
While the imagination of the majority of researchers we spoke with was shaped by the current capabilities of ChatGPT, the most technologically savvy participants were currently leveraging LLMs to do more complex tasks including systematic qualitative coding. Several participants wanted to use LLMs to do qualitative coding, but did not know how to use current tools, indicating their lack of use now could be partly due to a skills gap. Thus, the deployment of intentionally designed tools like CollabCoder (Gao et al., 2023a), SenseMate (Overney et al., 2024), and TIQA (Tseng et al., 2025) is even more important, since the state of accessible tooling may be actively limiting responsible research. We also see evidence that this divide in adoption could, on the other hand, be more fundamental: some participants simply deemed AI tools for tasks like coding to be needless departures from their qualitative sensemaking process.
5.3. Fundamental Tensions in Using LLMs for Qualitative Work
@@
A set of potentially fundamental tensions arose in considering the use of LLMs to conduct qualitative research. There has long been a tension between work seeking to combine computational power with interpretivist approaches (Baumer et al., 2017; Nelson, 2020) and work arguing that qualitative values must be protected from quantitative evaluation standards (Crabtree, 2024; Soden et al., 2024). Our interviews suggest that LLMs may continue blurring the line between positivist and interpretivist approaches, as exemplified in emerging work (Lam et al., 2024; Choksi et al., 2024). As Messeri and Crockett (2024) discuss, the way some approaches are more readily aided by LLMs may change the research landscape by making some questions more tractable than others. There is a danger that the ability to analyze data at scale with less human involvement may make situated, slow, contextual and interpretivist approaches to qualitative data less attractive, potentially because of incentives to quickly publish, or because of the perceived objectivity of LLMs compared to human analysts. Mixed-method researchers, who make up roughly half of our participants, may be particularly positioned to accelerate this change. Published work uses LLMs for many types of text analysis, including in powerful new tools that are increasingly positioned to facilitate “theory-driven analysis” of textual data (Lam et al., 2024), and these tools may be adopted by users from different epistemological traditions. As LLMs produce summaries, topics, thematic interpretations, and labels of textual data that look increasingly fluent, nuanced, and contextual, researchers may increasingly be tempted to use these outputs in analysis while decreasing the type of engagement with participants and their data that researchers say is central to qualitative research values.
6. Implications for Researchers and Designers
6.1. Considerations for researchers using LLMs in qualitative work
Thus far, we have explored how the combination of opaque tooling, lagging norms, and social context lead to an uncertain environment for qualitative researchers today. In this section, we answer RQ3 by providing an overview of key considerations and recommendations for LLM-curious qualitative researchers, or those who feel that LLM tools may interfere with their processes. Maintaining researcher agency and educating researchers on AI tools is vital as LLMs are embedded across more surfaces. We used prior work and our own expertise to reflect on the tensions raised by participants; we thus constructed Table LABEL:tab:choices to distill some main choices researchers can make relating to the uses and potential implications of LLMs in qualitative contexts.
Table 2. Table highlights the main considerations and recommendations regarding how qualitative researchers can incorporate LLMs into their research process. Considerations include a brief overview of the main tradeoffs involved in specific decisions, while Recommendations provide explicit guidelines for researchers that minimize risk to privacy, sensemaking, and research ethics.
Category Considerations Recommendations
Policies
Institutional, venue, and government guidelines and regulations Policies and regulations can help to provide basic guidance, but may be out of date or insufficient by themselves. Familiarize yourself with and follow institutional, publication, and federal policies, but treat these as a starting point for ethical engagement.
Research Ethics & Values
Consent Participants may not know that their data could be shared with a third-party company, or be aware of potential implications, such as the implication of models training on their data. Update consent forms on details if planning to use LLMs in the research, particularly in a training capacity, and educate participants on potential risks.
Privacy Guidelines may not strictly disallow using participant data with proprietary LLMs, even ones that train on inputs, leading to identification risk in the future. Additionally, LLMs may be embedded in the tools you currently use. Do not use proprietary models that train on input when interacting with participant data. Host local models and use open source models when possible. Anonymize confidential participant data if using a proprietary model. Confirm data policies before using a tool like MAXQDA that integrates LLM features.
Disclosure Venues may mandate specific forms of disclosure, but researchers may wish to disclose additional uses of LLMs for research. Reflexively engage with how LLMs may have shaped the relationship with participants, ideas in the project, and perceptions of the data. Proactively disclose LLM uses that significantly impacted a sensemaking process or anaysis outputs, including in ideation. When LLMs are used in analysis, discuss validation, including, disclosing model, model version, prompts, and performance, if appropriate.
Intentional task selection Chat-based interactions make unintentionally transitioning from one task to another easy. Tasks carry heterogeneous benefits and risks for qualitative research, and some decisions about LLM involvement or automation depend on research values. Interpretivist qualitative research tasks may be particularly at odds with LLM use. Before using an LLM, outline all tasks under consideration for LLM use. For each one, decide if the task is appropriate for LLM involvement or automation based on risks to compliance, originality, bias, privacy, and validity according to research values.
Validation LLMs should always be validated in task-specific ways in research. A clear way to validate the performance of an LLM may or may not exist for every task, and some tasks may be inappropriate for LLMs to perform even if they can be validated. Formal validation may be more tractable for positivist research, and a plan for validation should be made prior to engagement. Interpretivist and other research stances should consider whether any tasks are appropriate, given challenges to validation. Decide how to validate an LLM’s performance on any chosen task. This could include side by side comparison to human performance in qualitative coding, for example, or manual review of any edits to text suggested by an LLM. Only use an LLM if performance can be appropriately validated for the task.
LLM bias Consider LLMs’ cultural and political bias as you decide which tasks are appropriate for an LLM. Consider whether participant populations are aligned with LLM training data, and what the impacts of misalignments might be. For studies on populations unlikely to be well-represented in training data of off-the-shelf LLMs, avoid using off-the-shelf models in analysis to limit misinterpretation, misrepresentation, or hallucination, or consider fine-tuning.
Tool Selection
LLM chat interfaces Chat-based interactions are highly interactive and can support brainstorming in a research process, but interactivity can induce slippage across tasks and be difficult to validate for every use case. As such, some researchers may decide the risks of task-agnostic chat interfaces are not worthwhile. If using a chat interface, clearly outline the tasks for which the chat-based LLM will be used, and make a plan for validation before approaching the task. Monitor regularly for slippage across tasks.
QDA Software Software for QDA is intentionally designed for a qualitative coding use case, so it may provide validation (like measuring IRR across collaborators) or be designed with risks to sensemaking in mind. However, highly structured tools may miss opportunities for interactivity afforded by LLMs, so appropriateness depends on the intended task. Using structured tools reduce some but not all risks. If using LLM integrations into existing or emerging software tools, investigate the platform’s privacy policies and still consider the bias an LLM could introduce into your analysis.
Task-specific tools (e.g. Microsoft Excel, Grammarly) LLMs are increasingly embedded across the tools we use every day, even for simple tasks. Tools like Grammarly with more prescribed functionality constrained to editing may limit interference with your sensemaking process compared to open-ended tools. Reflect on your full set of research tools. Monitor for potential compromise of participant data if you use Grammarly or other tools to process or edit participant data.
Non chat-based LLM use Researchers may wish to use or create a research infrastructure using an LLM, either directly or accessed through an API, for tasks like procedurally annotating data. In these cases, choices between proprietary models and open source models, or locally hosted and remotely accessed models become salient, and carry different implications for privacy, transparency, and quality. If possible, use an open source model rather than a proprietary model to promote transparency and open science. If possible, use a locally hosted model for greater control over data. If using a proprietary model, anonymize data and do not use models or API services that train on input.
We now explain the three main covered areas of Table LABEL:tab:choices: Policies, Research Values & Ethics, and Tool Selection.
6.1.1. Policies
The first dimension in the table is the role of policies. We recommend actively keeping up to date on all policies from relevant communities, but treating these policies as an entry point rather than as a final answer on appropriate engagement with these tools. Policies may vary between venues; for example, some venues mandate specific types of disclosure (ACM, 2024) while others disallow the use of AI in some contexts (Sage, 2024).
As surfaced by our participants, existing policies and regulations may not engage deeply enough with the risks to certain participant populations. Therefore, while policies are important, they are likely insufficient considerations by themselves. The rest of the table provides a scaffold for considering the conscientious use of LLMs in qualitative work beyond established policies.
6.1.2. Research ethics and values
The second main dimension of the table encompasses Consent, Privacy, Disclosure, Intentional task selection, and Validation. Consent and privacy protection are fundamental values in qualitative research. Our participants indicated a current lack of scaffolding in current tools guiding how to use LLMs while preserving these values. Researchers should update their consent forms to accurately reflect intended uses of LLMs in research, as recommended by Kapania et al. (2024). These forms should include details on how participant privacy will be protected and education about potential risks.
Participants repeatedly highlighted a fear of compromising participant privacy. Researchers must take care not to share personally identifiable information or compromise the data of at-risk populations. These risks are particularly salient when using proprietary LLMs for any task involving confidential participant data. If using a proprietary model for such a task, carefully consider how to anonymize data, such as removing names or identifiable terms (e.g. (Choksi et al., 2024)). Avoid LLM interfaces which use your participants’ data as training data for the model. Check whether any tools with LLM-based features (e.g. Grammarly or Atlas.ti) train on inputs, and how they handle data privacy before using them. Privacy may be better protected by hosting open source models locally, but this is not currently feasible for all researchers.
Choosing which types of disclosure are appropriate regarding AI use is an opportunity for qualitative reflexivity. Qualitative researchers, as reflected in our interviews, already center reflexivity and positionality in their work, and can set precedent by explicitly engaging with how an LLM shaped a project. Researchers could consider the potential benefits of additional AI disclosure beyond those required by a venue, institution, or community. However, disclosing the use of AI may risk stigmatization, especially in non-computing fields (Liao et al., 2024), which should be balanced with the benefit additional transparency and context disclosure about AI usage could bring. Disclosures should ideally detail which models were used, and any fine-tuning, optimization, or prompt engineering performed.
Our table also outlines the value of intentional task selection when using LLMs in qualitative research. Intentional deployment of AI is key to maintaining researcher agency (Jiang et al., 2021; Feuston and Brubaker, 2021), but our participants highlight how general purpose tools now blur boundaries between tasks. With this challenge in mind, intentional task selection is even more pressing. The appropriateness of involving an LLM in a task can depend not only on venue policies, but also on research values, like an interpretivist or positivist stance. Before using an LLM for any task in your research, researchers should outline all tasks under consideration for LLM involvement and determine whether involving an LLM is appropriate based on their research values. In particular, researchers should consider how involving an LLM in the task may shape their perspective, ideas, and findings if used in ideation, analysis, or writing. Researchers should also consider whether they may be committing accidental acts of academic plagiarism by relying on LLM-generated outputs (Bhardwaz and Kumar, 2023) across the research pipeline, and should take responsibility for checking for copied content. If any tasks are determined to be appropriate for LLM involvement, researchers should monitor whether the planned task matches ongoing use of the LLM.
Our table also centers validation as a key consideration when using LLMs for qualitative research. Whether and how LLM usage can be validated appropriately depends on the task under consideration and individual research values. For example, if using an LLM to edit text, an appropriate validation might be to personally review each LLM-generated edit. If using an LLM to scale data annotation beyond possibility of human review, it may become appropriate or even necessary to use more positivist measures of validation, like inter-rater reliability, to measure accuracy compared to a human baseline. Not all qualitative research tasks have a clear appropriate validation: in particular, interpretivist researchers may find LLMs cannot be validated for tasks of interest and should be avoided. We advise making plan for validation of LLM performance for any chosen task before starting to interact with the tool, since anchoring bias can cause one to rationalize findings and outputs after already interacting with the LLM.
Additionally, when considering LLM appropriateness for any task, we recommend considering LLM training data, current performance, and biases. For example, LLMs may be useful for editing English text written by non-native English speakers, which appears to be a primary emerging use case from our participants and in other research (Liao et al., 2024). However, LLMs may not yet be advanced enough in other languages to be used for the same purpose. Model performance on tasks like transcription and analysis often directly reflects training data (Koenecke et al., 2020), so alignment between participant interests and a model’s training data should be key considerations. Using a mainstream LLM to analyze themes in a study of non-Western, non-English speaking participants may be inappropriate, since mainstream LLMs likely have poorer representations of those groups and viewpoints and could cause misinterpretation or harm. If LLM capabilities and bias cannot be considered and confidently addressed for a potential task, the tool should not be used.
6.1.3. Tool Selection
In the table, we outline the four types of LLM-aided tools that came up in interviews: LLM chat interfaces, QDA software, task-specific tools, and non-chat based LLM use.
Participants largely reported experiences with ChatGPT, which has interactivity and flexibility researchers enjoy but does not scaffold appropriate use for research purposes. Conversely, structured software solutions like MAXQDA for data analysis or Grammarly for text editing have LLM-based features but are designed to support users in a specific task-based use. Researchers should consider if a dedicated tool exists for their use case that guards against some risks to the user that open-ended chat interfaces pose. Chat is a useful way to interact with LLMs, but researchers may instead wish to integrate LLMs into software pipelines to procedurally generate outputs like data annotations.
The way an LLM is hosted and accessed also has implications for privacy and transparency. Proprietary models from companies like OpenAI can be accessed through a web browser or API, meaning data shared through those access points is shared with the company. Models from AI companies are often also closed source, meaning the model’s training data and weights are private, and understanding model behavior can be difficult. If using a proprietary model for research purposes, consider providing explicit justification as recommended by  Palmer et al. (2024). Open source models with known model weights and transparent training data are preferable for transparency, data privacy, and control, but currently, hosting them locally can be difficult and time-consuming without resources and technical skills.
We have identified the urgent need for transparent tools that center qualitative research use cases. Based on our findings from qualitative researchers interested in engaging with LLMs, tool builders could emphasize 1) interactivity, 2) data exploration, and 3) support for qualitative coding in future qualitative research tools. Participants consistently mentioned the useful nature of the interactivity LLMs afford through a chat-based interface: they could ideate, obtain feedback, and potentially make requests for data retrieval or formatting in natural language (e.g. (Izacard and Grave, 2021; Gao et al., 2023b)). There is a gap in tools that facilitate these use cases in addition to other, more structured uses. Second, LLMs could help researchers see their data in a new way, ideally widening rather than homogenizing the nature of their insights. Lastly, qualitative researchers are still interested in using LLMs to conscientiously assist in the coding process. Below are principles for tool designers when considering how to create tools that responsibly address these opportunity areas.
6.2.1. Design for participant privacy
Proprietary LLMs have sparked concerns regarding participant privacy. Tools that incorporate LLMs should be explicit about if, when, and how they call external APIs, and whether LLMs used are proprietary or open source. They should also explain to users before data is uploaded how they handle privacy, how they expect users to deanonymize data before using tools, how they scrub or plan to delete data, and whether they use inputs for training models. Second, some researchers mention open source models as a potential solution to privacy concerns, but many are unable to implement solutions to this problem themselves. As such, there is an opportunity area to support researchers in having greater control and transparency over data privacy. More tools that make it easy to locally host, steer, and fine-tune open source models for qualitative research functions could empower researchers.
6.2.2. Design for intentional use
The lines between structured tools built for a single task and open-ended chat-based tools are blurring. Chat-based LLMs allow flexibility and interactivity, but do not currently support intentional use for specific tasks in the research process, making slippage into unintentional uses of LLMs easy. Researchers want to maintain agency over if, when, and how they are assisted or influenced by AI tools in their research process. Tools could be designed to guide users to intentionally select appropriate uses of LLM-based tools, and also offer features that allow researchers agency to develop their own ideas before opting into LLM assistance in any task.
6.2.3. Design for transparency and validation
Tools should give researchers ways to transparently evaluate performance. Appropriate validation varies depending on research values, and on the specific research task. For brainstorming, designing for validation could involve designing features to understand or probe outcomes and suggestions. For positivist qualitative coding tasks, features designing for reproducibility of outcomes, interrater reliability, and error analysis from LLM annotation compared to human coding may be useful. Features that require researchers to review outputs may help accomplish this goal, depending on the task.
6.2.4. Design for researcher context
Researchers often wish to bring in their own context when performing qualitative analysis, including their past experiences, the texts that influence their work, and the theories in which they ground their work. Tools could embed unique researcher perspectives with models that are trained or fine-tuned according to particular data sets or perspectives of interest to the researcher.
6.2.5. Design for deep engagement with data
Qualitative researchers worry that LLM-based tools could distance them from data rather than deepen their relationship with it, but tools can be explicitly designed to foster ‘closeness’ with data (Gilbert, 2002). Many qualitative researchers expressed interest in using LLMs to see their data through another lens, but found that insights and themes produced on their data are generic or repetitive. Tools can explore LLMs as a way to produce more unique ways of understanding data, drawing researchers’ attention to less predictable features of the data, as suggested by Jiang et al. (2021), or helping researchers examine or challenge their existing theories with direct evidence from the data. These features may even deepen researchers’ understanding of the data.
6.2.6. Design to consider participant perspectives and interests
Researchers are concerned about the variable performance LLMs have across contexts, knowledge domains, cultures, and languages. Tools should work to facilitate better alignment between LLMs and diverse subject populations. When possible, tool builders could give the option to use models that better reflect a target, like models that specialize in a particular language or have specialized training data, and facilitate investigating where there might be errors or bias.